OpenStudy (anonymous):

HELP!!!! let f(x) = sqrt of -1-x if x<2 1 if x = -2 3x+8 if x>2 Calculate the limits: lim x of -2- f(x)=? lim x of -2+ f(x)=? lim x of -2 f(x)=?

OpenStudy (anonymous):

plug in -2 and see what you get

OpenStudy (anonymous):

i am assuming you meant \[f(x) = \left\{\begin{array}{rcc} \sqrt{-1-x} & \text{if} & x <-2 \\ 1& \text{if} & x=-2 \\ 3x+8 & \text{if} &x>-2 \end{array} \right.\]

OpenStudy (anonymous):

otherwise question doesn't really make sense

OpenStudy (zarkon):

showing off your piecewise function skills again I see :)

OpenStudy (anonymous):

lovingly stolen from zarkon (tm)

OpenStudy (anonymous):

actually i only stole the text part i had the rest from google

OpenStudy (anonymous):

thats correct except the top should be square root of -1-x +1

OpenStudy (anonymous):

like this? \[f(x) = \left\{\begin{array}{rcc} \sqrt{-1-x} +1& \text{if} & x <-2 \\ 1& \text{if} & x=-2 \\ 3x+8 & \text{if} &x>-2 \end{array} \right.\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in this case if you plug in -2 in the first formula you get 2

OpenStudy (anonymous):

and if you plug in -2 in the last formula you also get 2

OpenStudy (anonymous):

so the limit from the left is 2, the limit from the right is 2, and since they agree, that limit is 2

OpenStudy (anonymous):

ok so for these you just plug in for x?

OpenStudy (anonymous):

what the actual values is at x = -2 is irrelevant as far as the limit is concerned

OpenStudy (anonymous):

so that answer would be does not exist?

OpenStudy (anonymous):

yes don't make it as hard as your calc teacher makes it seem

OpenStudy (anonymous):

oh no that answer is the limit is 2

OpenStudy (anonymous):

let me write it out

OpenStudy (anonymous):

\[\lim_{x\rightarrow -2^-}=\sqrt{-1-(-2)}+1=2\] and \[\lim_{x\rightarrow -2^+}=3\times (-2)=8=2\] and since \[\lim_{x\rightarrow -2^-}f(x)=\lim_{x\rightarrow -2^+}f(x)=2\] then \[\lim_{x\rightarrow -2}f(x)=2\]

OpenStudy (anonymous):

typo line 3 should have been \[\lim_{x\rightarrow -2^+}=3\times (-2)+8=2\]

OpenStudy (anonymous):

ok thanx! what about this one: find in terms of the constant a. lim h of 0 5(a+h)squared - (5a)squared ____________________________ h

OpenStudy (anonymous):

is that \ \[\lim_{h \rightarrow 0}\frac{5(a+h)^2-5a^2}{h}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

two ther people on here gave me answers that were wrong. so im hoping you can help

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

my answers have been wrong also. im supposed to find in terms of the constant a.

OpenStudy (anonymous):

lets go slow. \[(a+h)^2=a^2+2ah+h^2\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[5(a+h)^2=5a^2+10ah+5h^2\]

OpenStudy (anonymous):

and so \[5(a+h)^2-5a^2=10ah+5h^2\]

OpenStudy (anonymous):

divide by h to get \[\frac{10ah+5h^2}{h}=\frac{h(10a+5h)}{h}=10a+5h\]

OpenStudy (anonymous):

now take the limit as h goes to zero get \[\lim_{h \rightarrow 0}\frac{5(a+h)^2-5a^2}{h}=\lim_{h \rightarrow 0}10a+5h=10a+5\times 0=10a\]

OpenStudy (anonymous):

YES! THANX!

OpenStudy (anonymous):

yw and don't fret, in a week you will do this in your head in 1 second

OpenStudy (anonymous):

you will say to yourself "the derivative of \[x^2\] is \[2x\] so the derivative of \[5x^2\] is \[10x\] finished

OpenStudy (anonymous):

AWESOME! I got some other questions Ill get back on later and see if we came up with the same answers!